how long will it take to $200 to triple itself if it is invested at 2% simple interest?
Answer by stanbon(48510)

how long will it take to $200 to triple itself if it is invested at 2% simple interest?
-------------------------------------
I = Prt
A(t) = P + Prt
A(t) = P(1+rt)
-----------------------
Your Problem:
3*200 = 200(1+0.02*t)
3 = 1 + 0.02t
2 = 0.02t
t = 2/(0.02)
t = 2/(2/100)
t = 100 years
=======================
Cheers,
Stan H.